OpenStudy (anonymous):

chapter 2:Quadratic Equations

OpenStudy (anonymous):

If a and B are the roots of the quadratic equation \[\frac{ 2x-3 }{ 8 }=\frac{ 3-2x }{ 5x },\]find the value of \[a^2+b^2\]

OpenStudy (welshfella):

the standard form for a quadratic equation is ax^2 + bx + c = 0 product of roots AB = c/a and A + B = -b/a first rearrange your equation to the standard form

OpenStudy (anonymous):

i got \[10x^2+x-24=0\]

OpenStudy (welshfella):

then to find A^2 + B^2 use the identity A^2 + B^2 = (A + B)^2 - 2AB

OpenStudy (anonymous):

\[a+B=-\frac{ 1 }{ 10 }=SOR\]\[aB=-\frac{ 12 }{ 5 }=POR\]

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

I don't understand how u get -2aB

OpenStudy (welshfella):

(a +B)^2 = a^2 +2aB + B^2

OpenStudy (welshfella):

now take away the 2aB

OpenStudy (anonymous):

Okay

OpenStudy (welshfella):

now you just plug in the values for a +B and aB

OpenStudy (anonymous):

\[(a+B)^2-2aB=(-\frac{ 1 }{ 10 })^2-2(-\frac{ 12 }{ 5 })\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ 100 }+\frac{ 24 }{ 5 }\]

OpenStudy (anonymous):

\[=\frac{ 481 }{ 100 }\]

OpenStudy (anonymous):

Thnx @welshfella

OpenStudy (welshfella):

looks good yw